At an amusement park there is a ride in which cylindrically shaped chambers spin around a central axis. People sit in seats facing the axis, their backs against the outer wall. At one instant the outer wall moves at a speed of 3.0 m/s, and an 83-kg person feels a 600-N force pressing against his back. What is the radius of the chamber?
Answer #1

Centripetal force, F = mv^2/r

Radius, r = mv^2/F = (83 x 3^2)/600 = 1.245 m

Comment in case any doubt please rate my answer.. ..